# Monotone Neural Barrier Certificates

### v1

@@ 2.2 Monotone Dynamical Systems
Definition 4 (Monotone maps).
@@
Definition 5 (Monotone dynamical systems).
@@
For a differentiable transition map $f$ and convex sets $\mathcal{X}$ and $U$, the system $\mathfrak{S}$ is monotone if and only if $Df(x,u)\geq 0_{n\times(n+m)}$ for all $(x,u)\in\mathcal{X}\times U$ HLS:95 . Monotone systems preserve the standard partial order along their trajectories HLS:95 ; angeli2003monotone , resulting in highly regular behaviors that facilitate reasoning about dynamics via partial orders.
@@
We consider the safety of a monotone dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$ with an unknown transition map $f$. While $f$ is not explicitly known, we assume access to a simulator that returns $f(x,u)$ for any $x\in\mathcal{X}$ and $u\in U$. This paper focuses on the following problem.
Problem 1 (Safety Certificates for Unknown Monotone Systems).
@@
The monotonicity assumption in Problem 1, although made in the absence of an explicit model, is often justified by prior knowledge of the system’s physical properties. Many physical processes naturally exhibit monotonicity, where increases in inputs or states lead to non-decreasing responses. This property is common in domains such as traffic networks SC-MA:15 , power systems park2020uniqueness , biological processes EDS:07 , and supply chains yu2019equilibrium .
3 Monotone Neural Barrier Certificates
We propose a data-driven framework that leverages system monotonicity to learn barrier functions using neural networks. This approach enables the construction of sound safety certificates from limited data, avoiding full system identification or exhaustive sampling. We first analyze the structure of barrier certificates for monotone systems and show that, under certain assumptions on the unsafe set $\mathcal{X}_{u}$, any safe monotone system admits a monotone robust (resp. control) barrier certificate.
3.1 Monotone Barrier Certificates
In this subsection, we analyze the structure of barrier certificates for safety verification and control synthesis in a monotone system $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$. We show that if the unsafe set $\mathcal{X}_{u}$ is upper closed, both problems can be addressed using monotone barrier certificates.
Theorem 3 (Monotone Barrier Certificates).
Consider a monotone dtDs $\mathfrak{S}:=(\mathcal{X},\mathcal{X}_{0},U,f)$ with an upper-closed set of unsafe states $\mathcal{X}_{u}\subseteq\mathcal{X}$. Then, the following statements hold:
  1. 1.
    (Safety verification). $\mathfrak{S}$ is safe with respect to $\mathcal{X}_{u}$, for every input $\{u_{t}\}_{t=0}^{\infty}\in U$ if and only if there exists a monotone robust barrier certificate satisfying (2),(3), and (4).
  2. 2.
    (Safe control synthesis). there exists a map $K:\mathcal{X}\to U$ such that $\mathfrak{S}$ with the feedback controller $\{u_{t}=K(x_{t})\}_{t=0}^{\infty}$ is safe with respect to $\mathcal{X}_{u}$ if and only if there exists a monotone control barrier certificate satisfying (2),(3), and (5).
Proof. We refer to Appendix D for the proof.
For monotone systems, Theorem 3 reduces the search for robust (control) barrier certificates to monotone barrier functions. As will be shown in Section 3.2, the monotonicity structure of barrier certificates plays a crucial role in our framework, leading to significantly lower sample complexity and computational cost—particularly in high-dimensional systems. This structure also enables restricting learning-based approximators to inherently monotone functions. Monotone neural networks thus provide a natural and expressive choice for approximating such certificates.
@@ 3.1 Monotone Barrier Certificates
Let $y_{i-1}$ and $y_{i}$ denote the input and output of the $i$-th layer, with activation function $\sigma:\to$. The network $N$ is monotone if (i) all weight matrices satisfy $W_{i}\geq 0_{n_{i}\times n_{i+1}}$ and (ii) $\sigma$ is non-decreasing. Monotone neural networks are universal approximators of monotone functions daniels2010monotone .
Using Theorem 3, the search for robust (resp. control) barrier certificates reduces to finding monotone neural networks satisfying conditions (2), (3), and (4) (resp. (5)). However, formally verifying such certificates remains intractable due to the need to evaluate the network over a continuum of states. In the next section, we show how the system’s and certificate’s monotonicity enable safety verification and control synthesis from finite data.
3.2 Monotone Barrier Certificates from Finite Data
@@
$\mathcal{X}_{0}\subseteq\bigcup_{i\in\mathcal{I}}[\underline{x}^{i},\overline{x}^{i}],\qquad\mathcal{X}_{u}\subseteq\bigcup_{i\in\mathcal{U}}[\underline{x}^{i},\overline{x}^{i}].$ (6)
@@
By construction, any $x\in\mathcal{X}_{0}$ satisfies $x\leq\overline{x}^{i}$ for some $i\in\mathcal{I}$, and any $x\in\mathcal{X}_{u}$ satisfies $\underline{x}^{i}\leq x$ for some $i\in\mathcal{U}$. For monotone systems with monotone barrier certificates, the points $\{\underline{x}^{i}\}_{i\in\mathcal{U}}$ and $\{\overline{x}^{i}\}_{i\in\mathcal{I}}$ act as worst-case representatives. Thus, the search for robust barrier certificates satisfying (2)-(4) can be conservatively reformulated as finding a monotone neural network $N:\mathcal{X}\to\mathbb{R}$ satisfying:
$N(\overline{x}^{i})$ $\leq\gamma\quad$ $\forall i\in\mathcal{I},$ (7)
$N(\underline{x}^{i})$ $>\eta\quad$ $\forall i\in\mathcal{U},$ (8)
$N(\underline{x}^{i})$ $\leq\delta\quad\implies\quad\max_{j\in\mathcal{S}}N(f(\overline{x}^{i},\overline{u}^{j}))\leq\delta\quad$ $\forall i\in\mathcal{P}.$ (9)
@@
A monotone neural network $N:\mathcal{X}\to$ satisfying (7), (8), and (9) is called a monotone neural robust barrier certificate. Similarly, for safe control synthesis, the search can be reformulated as finding a monotone neural network $N:\mathcal{X}\to\mathbb{R}$ satisfying (7), (8), and
$N(\underline{x}^{i})$ $\leq\delta\quad\implies\quad\min_{j\in\mathcal{S}}N(f(\overline{x}^{i},\underline{u}^{j}))\leq\delta\quad$ $\forall i\in\mathcal{P}.$ (10)
@@
A monotone neural network $N:\mathcal{X}\to$ that satisfies conditions (7), (8), and (9) is referred to as a monotone neural control barrier certificate. Conditions (7)–(10) reduce the verification of neural barrier certificates from the entire (continuous) state set to a finite set of carefully selected maximal and minimal points derived from the coverings of $\mathcal{X}$ and $U$. The following result provides formal guarantees for safety verification and safe control synthesis using monotone neural barrier certificates.
Theorem 4 (Monotone Neural Barrier Certificates).
@@
Consider a monotone dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$ with an unknown state transition map $f$ and a set of unsafe states $\mathcal{X}_{u}$. Let $\{[\underline{x}^{i},\overline{x}^{i}]\}_{i\in\mathcal{P}}$ be a family of hyper-rectangles that cover the state set $\mathcal{X}$, $\{[\underline{u}^{i},\overline{u}^{i}]\}_{i\in\mathcal{S}}$ be a family of hyper-rectangles that cover the input set $U$, and the index sets $\mathcal{I},\mathcal{U}$ be the minimal subsets of $\mathcal{P}$ satisfying (6). Then,
  1. 1.
@@
    (Safety verification) If there exists a monotone neural network $N:\mathcal{X}\to$ satisfying (7), (8), and (9), with $\gamma\leq\delta\leq\eta$, then $\mathfrak{S}$ is safe with respect to $\mathcal{X}_{u}$, for every input $\{u_{t}\}_{t=0}^{\infty}\in U$;
  2. 2.
@@
    (Safe control synthesis) If there exists a monotone neural network $N:\mathcal{X}\to$ satisfying (7), (8), and (10), with $\gamma\leq\delta\leq\eta$, then there exists a map $K:\mathcal{X}\to U$ such that $\mathfrak{S}$ with the feedback controller $\{u_{t}=K(x_{t})\}_{t=0}^{\infty}$ is safe with respect to $\mathcal{X}_{u}$.
Proof. We refer to Appendix E for the proof.
@@
Theorem 4 highlights key advantages of our data-driven approach. It allows constructing safety certificates from arbitrary state space partitions, avoiding the curse of dimensionality in fine discretizations anand2023formally and the reliance on SMT solvers for neural network verification edwards2024fossil . By leveraging system monotonicity, it scales to high-dimensional systems and does not require knowledge of Lipschitz constants for the dynamics or barrier function.
3.3 Training Monotone Barrier Certificates.
@@
In this section, we present an adaptive algorithm for training monotone neural robust (resp. control) barrier certificates. Let $\{[\underline{x}^{i},\overline{x}^{i}]\}_{i\in\mathcal{P}}$ and $\{[\underline{u}^{i},\overline{u}^{i}]\}_{i\in\mathcal{S}}$ be hyper-rectangular coverings of the state and input spaces, respectively, and $\mathcal{I},\mathcal{U}$ the index sets defined in (6).
@@
$L_{1}$ $:=\mathrm{MSE}(N(\overline{x}^{i}),\gamma),$ $\text{for all }i\in\mathcal{I},$ (11)
$L_{2}$ $:=\mathrm{MSE}(N(\underline{x}^{i}),\eta),$ $\text{for all }i\in\mathcal{U},$ (12)
$L_{3}$ $:=\mathrm{MSE}\left(\min_{j\in\mathcal{S}}N(f(\overline{x}^{i},\underline{u}^{j})),\delta\right),$ $\text{for all }i\in\mathcal{P}.$ (13)
Training uses gradient-based optimization on $L_{\mathrm{total}}$, with tunable parameters $\gamma,\delta,\eta\in\mathbb{R}_{>0}$ satisfying $\gamma<\delta<\eta$. The process stops once $N$ meets conditions (7), (8), and (9).
For safe control synthesis, we jointly train two monotone neural networks: $N:\mathcal{X}\to\mathbb{R}$, which serves as the control barrier certificate, and $K:\mathcal{X}\to\mathcal{U}$, which represents the feedback policy. The objective for both of these networks is the composite mean-squared-error loss $\widehat{L}_{\mathrm{total}}:=L_{1}+L_{2}+L_{4}$, where $L_{1}$ and $L_{2}$ are defined in (11) and (12), and $L_{4}$ is given by
$L_{4}$ $:=\mathrm{MSE}\left(N(f(\overline{x}^{i},K(\overline{x}^{i}))),\delta\right),$ $\forall i\in\mathcal{P}.$ (14)
Training is performed jointly on $N$ and $K$ using gradient-based optimization on $\widehat{L}_{\mathrm{total}}$, with tunable parameters $\gamma,\delta,\eta\in\mathbb{R}_{>0}$ such that $\gamma\leq\delta\leq\eta$. It stops when $N$ satisfies (7), (8), and the following condition:

### v3

@@ 2.2. Monotone Dynamical Systems
Definition 2.6 (Monotone maps).
@@
Definition 2.7 (Monotone dynamical systems).
@@
For a differentiable transition map $f$ and convex sets $\mathcal{X}$ and $U$, the system $\mathfrak{S}$ is monotone if and only if $Df(x,u)\geq 0_{n\times(n+m)}$ for all $(x,u)\in\mathcal{X}\times U$ [29]. Monotone systems preserve the standard partial order along their trajectories [29, 5], resulting in highly regular behaviors that facilitate reasoning about dynamics via partial orders.
@@
We consider the safety of a monotone dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$ with an unknown transition map $f$. While $f$ is not explicitly known, we assume access to a simulator that returns $f(x,u)$ for any $x\in\mathcal{X}$ and $u\in U$. This report focuses on the following problem.
Problem 2.8 (Safety Certificates for Unknown Monotone Systems).
@@
The monotonicity assumption in Problem 2.8, although made in the absence of an explicit model, is often justified by prior knowledge of the system’s physical properties. Many physical processes naturally exhibit monotonicity, where increases in inputs or states lead to non-decreasing responses. This property is common in domains such as traffic networks [10], power systems [22], biological processes [30], and supply chains [33].
3. Monotone Neural Barrier Certificates
We propose a data-driven framework that leverages system monotonicity to learn barrier functions using neural networks. This approach enables the construction of sound safety certificates from limited data, avoiding full system identification or exhaustive sampling.
3.1. Monotone Barrier Certificates
In this subsection, we propose a suitable structure of barrier certificates that enables efficient data-driven safety verification and control synthesis for the monotone system $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$.
Theorem 3.1 (Monotone Barrier Certificates).
Consider a monotone dtDs $\mathfrak{S}:=(\mathcal{X},\mathcal{X}_{0},U,f)$ with an unsafe states $\mathcal{X}_{u}\subseteq\mathcal{X}$. Then, the following statements hold:
  1. (1)
    (Safety verification). $\mathfrak{S}$ is safe with respect to $\mathcal{X}_{u}$, for every input $\{u_{t}\}_{t=0}^{\infty}\in U$ if there exists a monotone robust barrier certificate satisfying (2.2),(2.3), and (2.4).
  2. (2)
    (Safe control synthesis). there exists a map $K:\mathcal{X}\to U$ such that $\mathfrak{S}$ with the feedback controller $\{u_{t}=K(x_{t})\}_{t=0}^{\infty}$ is safe with respect to $\mathcal{X}_{u}$ if there exists a monotone control barrier certificate satisfying (2.2),(2.3), and (2.5).
Proof. We refer to Appendix A for the proof.
For monotone systems, Theorem 3.1 provides a sufficient condition for searching for robust (control) barrier certificates within the class of monotone barrier functions. As will be shown in Section 3.2, the monotonicity structure of barrier certificates plays a crucial role in our framework, leading to significantly lower sample complexity and computational cost—particularly in high-dimensional systems. This structure also enables restricting learning-based approximators to inherently monotone functions. Monotone neural networks thus provide a natural and expressive choice for approximating such certificates.
@@ 3.1. Monotone Barrier Certificates
Let $y_{i-1}$ and $y_{i}$ denote the input and output of the $i$-th layer, with activation function $\sigma:\mathbb{R}\to\mathbb{R}$. The network $N$ is monotone if (i) all weight matrices satisfy $W_{i}\geq 0_{n_{i}\times n_{i+1}}$ and (ii) $\sigma$ is non-decreasing. Monotone neural networks are universal approximators of monotone functions [11].
By Theorem 3.1, the search for robust (respectively, control) barrier certificates reduces to finding monotone neural networks that satisfy conditions (2.2), (2.3), and (2.4) (respectively, (2.5)). However, formally verifying such certificates remains intractable due to the need to evaluate the network over a continuum of states. In the next section, we show how the system’s and certificate’s monotonicity enable safety verification and control synthesis from finite data.
3.2. Monotone Barrier Certificates from Finite Data
@@
(3.1) $\mathcal{X}_{0}\subseteq\bigcup_{i\in\mathcal{I}}[\underline{x}^{i},\overline{x}^{i}],\qquad\mathcal{X}_{u}\subseteq\bigcup_{i\in\mathcal{U}}[\underline{x}^{i},\overline{x}^{i}].$
@@
By construction, any $x\in\mathcal{X}_{0}$ satisfies $x\leq\overline{x}^{i}$ for some $i\in\mathcal{I}$, and any $x\in\mathcal{X}_{u}$ satisfies $\underline{x}^{i}\leq x$ for some $i\in\mathcal{U}$. For monotone systems with monotone barrier certificates, the points $\{\underline{x}^{i}\}_{i\in\mathcal{U}}$ and $\{\overline{x}^{i}\}_{i\in\mathcal{I}}$ act as worst-case representatives. Thus, the search for robust barrier certificates satisfying (2.2)-(2.4) can be conservatively reformulated as finding a monotone neural network $N:\mathcal{X}\to\mathbb{R}$ satisfying:
(3.2) $N(\overline{x}^{i})$ $\leq\gamma\quad$ $\forall i\in\mathcal{I},$
(3.3) $N(\underline{x}^{i})$ $>\eta\quad$ $\forall i\in\mathcal{U},$
(3.4) $N(\underline{x}^{i})$ $\leq\delta\quad\implies\quad\max_{j\in\mathcal{S}}N(f(\overline{x}^{i},\overline{u}^{j}))\leq\delta\quad$ $\forall i\in\mathcal{P}.$
@@
A monotone neural network $N:\mathcal{X}\to\mathbb{R}$ satisfying (3.2), (3.3), and (3.4) is called a monotone neural robust barrier certificate. Similarly, for safe control synthesis, the search can be reformulated as finding a monotone neural network $N:\mathcal{X}\to\mathbb{R}$ satisfying (3.2), (3.3), and
(3.5) $N(\underline{x}^{i})$ $\leq\delta\quad\implies\quad\min_{j\in\mathcal{S}}N(f(\overline{x}^{i},\underline{u}^{j}))\leq\delta\quad$ $\forall i\in\mathcal{P}.$
@@
A monotone neural network $N:\mathcal{X}\to\mathbb{R}$ that satisfies conditions (3.2), (3.3), and (3.4) is referred to as a monotone neural control barrier certificate. Conditions (3.2)–(3.5) reduce the verification of neural barrier certificates from the entire (continuous) state set to a finite set of carefully selected maximal and minimal points derived from the coverings of $\mathcal{X}$ and $U$. The following result provides formal guarantees for safety verification and safe control synthesis using monotone neural barrier certificates.
Theorem 3.2 (Monotone Neural Barrier Certificates).
@@
Consider a monotone dtDS $\mathfrak{S}=(\mathcal{X},\mathcal{X}_{0},U,f)$ with an unknown state transition map $f$ and a set of unsafe states $\mathcal{X}_{u}$. Let $\{[\underline{x}^{i},\overline{x}^{i}]\}_{i\in\mathcal{P}}$ be a family of hyper-rectangles that cover the state set $\mathcal{X}$, $\{[\underline{u}^{i},\overline{u}^{i}]\}_{i\in\mathcal{S}}$ be a family of hyper-rectangles that cover the input set $U$, and the index sets $\mathcal{I},\mathcal{U}$ be the minimal subsets of $\mathcal{P}$ satisfying (3.1). Then,
  1. (1)
@@
    (Safety verification) If there exists a monotone neural network $N:\mathcal{X}\to\mathbb{R}$ satisfying (3.2), (3.3), and (3.4), with $\gamma\leq\delta\leq\eta$, then $\mathfrak{S}$ is safe with respect to $\mathcal{X}_{u}$, for every input $\{u_{t}\}_{t=0}^{\infty}\in U$;
  2. (2)
@@
    (Safe control synthesis) If there exists a monotone neural network $N:\mathcal{X}\to\mathbb{R}$ satisfying (3.2), (3.3), and (3.5), with $\gamma\leq\delta\leq\eta$, then there exists a map $K:\mathcal{X}\to U$ such that $\mathfrak{S}$ with the feedback controller $\{u_{t}=K(x_{t})\}_{t=0}^{\infty}$ is safe with respect to $\mathcal{X}_{u}$.
Proof. We refer to Appendix B for the proof.
@@
Theorem 3.2 highlights key advantages of our data-driven approach. It allows constructing safety certificates from arbitrary state space partitions, avoiding the curse of dimensionality in fine discretizations [3] and the reliance on SMT solvers for neural network verification [14]. By leveraging system monotonicity, it scales to high-dimensional systems and does not require knowledge of Lipschitz constants for the dynamics or barrier function.
3.3. Training Monotone Barrier Certificates.
@@
In this section, we present an adaptive algorithm for training monotone neural robust (resp. control) barrier certificates. Let $\{[\underline{x}^{i},\overline{x}^{i}]\}_{i\in\mathcal{P}}$ and $\{[\underline{u}^{i},\overline{u}^{i}]\}_{i\in\mathcal{S}}$ be hyper-rectangular coverings of the state and input spaces, respectively, and $\mathcal{I},\mathcal{U}$ the index sets defined in (3.1).
@@
(3.6) $L_{1}$ $:=\mathrm{MSE}(N(\overline{x}^{i}),\gamma),$ $\text{for all }i\in\mathcal{I},$
(3.7) $L_{2}$ $:=\mathrm{MSE}(N(\underline{x}^{i}),\eta),$ $\text{for all }i\in\mathcal{U},$
(3.8) $L_{3}$ $:=\mathrm{MSE}\left(\min_{j\in\mathcal{S}}N(f(\overline{x}^{i},\underline{u}^{j})),\delta\right),$ $\text{for all }i\in\mathcal{P}.$
Training uses gradient-based optimization on $L_{\mathrm{total}}$, with tunable parameters $\gamma,\delta,\eta\in\mathbb{R}_{>0}$ satisfying $\gamma<\delta<\eta$. The process stops once $N$ meets conditions (3.2), (3.3), and (3.4).
For safe control synthesis, we jointly train two monotone neural networks: $N:\mathcal{X}\to\mathbb{R}$, which serves as the control barrier certificate, and $K:\mathcal{X}\to\mathcal{U}$, which represents the feedback policy. The objective for both of these networks is the composite mean-squared-error loss $\widehat{L}_{\mathrm{total}}:=L_{1}+L_{2}+L_{4}$, where $L_{1}$ and $L_{2}$ are defined in (3.6) and (3.7), and $L_{4}$ is given by
(3.9) $L_{4}$ $:=\mathrm{MSE}\left(N(f(\overline{x}^{i},K(\overline{x}^{i}))),\delta\right),$ $\mbox{ for all }i\in\mathcal{P}.$
Training is performed jointly on $N$ and $K$ using gradient-based optimization on $\widehat{L}_{\mathrm{total}}$, with tunable parameters $\gamma,\delta,\eta\in\mathbb{R}_{>0}$ such that $\gamma\leq\delta\leq\eta$. It stops when $N$ satisfies (3.2), (3.3), and the following condition: